OpenStudy (cutepuppies):

There are 4 elephants in the preserve. Each elephant eats 2 1/4 tons of vegetation in a week. How much vegetation do the elephants eat in a week? Express your answer in simplest form. tons pleas try your best on solving make sure you are correct

OpenStudy (jdoe0001):

4 elephants each eats 2 1/4 ton 4 * 2 1/4 :)

OpenStudy (iluvsoccer):

\[\frac{ 4 }{ 1 }\times \frac{ 9 }{ 4 }\]

OpenStudy (jdoe0001):

9/4? where did that come from?

OpenStudy (iluvsoccer):

We can cross out the 4 as they rule out.

OpenStudy (jdoe0001):

and 4/1? wassdat?

OpenStudy (jdoe0001):

anyway.... so, what do you get from 4 * 2 1/4?

OpenStudy (iluvsoccer):

Don't you have to do that in order to multiply fractions? :/ Convert them to improper fractions? :/

OpenStudy (iluvsoccer):

I haven't done these in a while. :P

OpenStudy (jdoe0001):

dunno, let's ask @cutepuppies , I think she'd know

OpenStudy (jdoe0001):

so, what do you get from 4 * 2 1/4 @cutepuppies ? *biting nails* *profuse sweating* don't leave us in suspense

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ a}}\frac{{\color{brown}{ b}}}{{\color{purple}{ c}}}\implies \cfrac{{\color{blue}{ a}}\times {\color{purple}{ c}}+{\color{brown}{ b}}}{{\color{purple}{ c}}}\impliedby \textit{to make the mixed, to improper fraction}\) so... do that with 2 1/4, see what you get and then just multiply away